# Integrating Multi-Modal Input Token Mixer into Mamba-Based Decision Models: Decision MetaMamba

### v2

@@ Introduction
The convolution module aims to integrate neighboring embeddings within the window size. It applies filters across channels of the hidden dimension using a 1D convolution kernel. As illustrated in Figure 4a, to account for multi-modality, three different convolution kernels are used. This results in three distinct hidden states, generated by each kernel in the order of Rtg, state, and action. These hidden states are then combined to form a new hidden state tensor. The output tensor size matches the input tensor size.
1D convolution cannot account for positional information between inputs. Although methods such as (Sabour, Frosst, and Hinton 2017) have been developed to address this issue, they are too computationally intensive for decision models. With a filter length of 6, only two time steps $(rtg_{t-1},state_{t-1},action_{t-1},rtg_{t},state_{t},action_{t})$are considered. Although information outside the convolution window can be integrated through layer stacking, it is challenging for the entire information to be encompassed within the final step hidden state.
@@
As can be observed in Figure 4b, Zero-padded sequence vectors are grouped into embedding sets of window size. To perform computation in a single linear transformation without iteration, each window size vector within the set is flattened into a 1-dimensional vector. These are then transformed back into vectors with the original hidden state dimension via a linear layer. While the convolution module applies filters to the hidden state dimension, the linear module concatenates consecutive sequence vectors of window size in the time domain in a flattened manner. The resulting vectors, transformed back to the hidden state dimension, form a sequence, yielding an output tensor that matches the input tensor size. The linear layer also uses three different modules to distinguish between each modality.
@@
In the train trajectory dataset, each sequence of length, which includes Rtg, state, and action, is embedded into vectors of the same size through an input embedding layer. This process addresses the issue arising from the differing sizes of these components. Traditional transformer architectures typically add positional encoding to the embeddings. However, DMM ensures input causality through padding and processes the inputs using a sequence model, thereby obviating the need for positional encoding. By eliminating additional information dependent on the trajectory, the likelihood of behavioral cloning is further reduced. The output from passing through $n$ MetaMamba blocks is transformed into an action-dimension vector, resulting in the predicted action. The loss is then computed between the predicted actions and the true actions across all sequence lengths. As described, the loss is calculated for all predicted actions within a sequence, but during evaluation, only the final action in the sequence length is used.
@@
While a longer sequence length during training can provide more guidance, it may inversely affect the accuracy of the final action. In a situation where there is a shortage of necessary guides for learning, DMM was able to enhance the accuracy of the final action by considering both short-range and long-range information within the sequence. Therefore, DMM utilized shorter sequence lengths in all environments compared to existing models.
@@
In the evaluation stage, DMM interacts with the environment in real-time to infer the action token. Therefore, it is necessary to set the initial Rtg. Consistent with (Chen et al. 2021)., we set the initial target return to be higher than the actual return of the trajectory.
@@ Conclusion
We propose Decision Metamamba, a sequential decision-making model that employs the selective scan State-Space Model with a modified input layer. The input sequence in offline reinforcement learning includes data in different formats: state, returns-to-go, and action. To handle this, we utilized two distinct token mixer networks, collectively referred to as the Multi-Modal Token Mixer. These two token mixers are, respectively, token mixers that utilize a 1D convolution layer operating in the hidden state dimension and a linear layer for integrating tokens along the sequence dimension. Each network demonstrated strong performance in specific experimental domains.
To appropriately consider information from both short sequences close to the current step and distant steps, we employed the selective scan SSM. Unlike previous sequence models, the Selective Scan SSM facilitates rapid learning through parallel computation while utilizing a relatively small number of parameters, as some layers perform convolutional operations. This combination allows the model to achieve scores comparable to or exceeding those of existing models, despite having a lower number of parameters. Moreover, leveraging the recurrent properties of the transition matrix allows for inference speeds up to five times faster than traditional transformer models. Additionally, since the output of the token mixer sequentially serves as the input to the SSM for action inference, the model does not require time-step encoding, contributing to its enhanced generalizability.
The ongoing advancements from the initial Structured SSM to Gated SSM and Recurrent SSM (Patro and Agneeswaran 2024a) suggest that this study can pave the way for improved performance in offline RL decision models through the use of advanced SSMs.

### v3

@@ Introduction
The convolution module aims to integrate neighboring embeddings within the window size. It applies filters across channels of the hidden dimension using a 1D convolution kernel. As illustrated in Figure 4a, to account for multi-modality, three different convolution kernels are used. This results in three distinct hidden states, generated by each kernel in the order of Rtg, state, and action. These hidden states are then combined one by one per filter to form a new hidden state tensor. The output tensor size matches the input tensor size.
1D convolution cannot account for positional information. Although methods such as (Sabour, Frosst, and Hinton 2017) have been developed to address this issue, they are too computationally intensive for decision models. However, when setting the filter length to 6, only two steps $(rtg_{t-1},state_{t-1},action_{t-1},rtg_{t},state_{t},action_{t})$are considered, so the order becomes less important. Simultaneously, the effect of token mixing, which preserves the information from nearby steps, is adequately achieved.
@@
As depicted in Figure 4b, Instead of applying zero-padding for 1D convolution in the previous token mixer, zero vectors are added to the front to maintain causality. Subsequently, the window-size number of hidden states are flattened into a 1D vector and fed into a linear layer. These are then transformed back into a vector with the original hidden state dimension via a linear layer. While the convolution module applies filters to the hidden dimension, the linear module concatenates consecutive vectors of window size in the time domain. The resulting vectors from a sequence, yielding an output tensor that matches the input tensor size. Three distinct linear layers are utilized to differentiate between each modality
@@
To train, the model processes input sequences through three Metamamba layers to predict the output action. Each input sequence consists of 8 steps, where each step is composed of a state, action, and $rtg$. Before being fed into the Metamamba layers, an embedding network transforms the state, action, and $rtg$, which have varying dimensions, into fixed-size vectors. Unlike Transformer architectures that typically encode position into embeddings, the proposed DMM approach eliminates the need for positional encoding. Instead, DMM leverages a sequence model to process the inputs and enforces causality via padding. This mitigates the inclusion of trajectory-dependent positional information, thereby reducing the likelihood of behavioral cloning.
@@
The output sequence produced by the Metamamba layers retains the same dimensionality as the input sequence. For each step in the sequence, the model computes the loss between the predicted action and the target action. While training with longer sequences can provide additional context for action prediction, it may negatively impact the accuracy of the final action in the sequence during evaluation. DMM achieves higher accuracy in both training and evaluation phases by utilizing shorter sequences compared to traditional Transformer-based models.
@@
During the evaluation, DMM interacts with the environment in real time to predict the next action. Upon receiving the reward and next state from the environment, these are appended to the input sequence, and the model infers the subsequent action from the updated sequence.
In the evaluation setting, where the $rtg$ is unknown, an arbitrary initial $rtg$ is assigned(Chen et al. 2021). The received reward is then subtracted from this initial value to compute the updated $rtg$ for the next step. To evaluate the model’s performance under varying conditions, we conducted experiments using four different initial target returns, each exceeding the actual return of the trajectory.
@@ Conclusion
In this paper, we proposed a novel decision model for offline reinforcement learning (RL) called Decision MetaMamba (DMM). DMM introduces a modification to the token mixer in Mamba’s input layer, enabling adjacent step tokens to be mixed before being processed by the selective scan SSM. By incorporating a residual sum connection at the output, DMM reduces information loss from selective scan SSM and achieves higher performance compared to existing Mamba-based models. Experimental results demonstrated that this performance improvement is robust, regardless of the specific design of the token mixer. Additionally, incorporating multi-modality into the token mixer further enhanced performance.
DMM outperformed existing Transformer-based models while maintaining the advantages of Mamba, such as parameter efficiency and fast inference speed. This highlights the potential of DMM to extend Mamba’s capabilities to offline RL tasks.
The integration of SSMs through the modified input layer represents a novel and versatile approach that can be easily applied to other SSM architectures. This opens new possibilities for performance improvements across various SSM models, from the original structured SSM to Gated SSM and Recurrent SSM(Patro and Agneeswaran 2024a).